What is the next number in series 1,3,7,25,103,?
 
(i) 526
(ii) 521
(iii) 515
(iv) 509

Grade:8

4 Answers

Shubham Kamleshbhai Patel
7 years ago
I feel that there is some mistake in the question since...there is no relation between the 4th and 3rd, 5th and 4th ..
 
Please recheck the question.....
Akshit Khanna
7 years ago
I too feel the same but this question has been given in FIITJEE mental ability book for Udaya-I for class 8. In book its answer has been given as 521 but no logic has been given for it. If anyone can relate any logic to it, it will be great.
Rudrajyoti Das
6 years ago
1 * 1 + 2 =3
3 * 2 + 1=7
7 * 3 + 4=25
25 * 4 + 3=103
103 * 5 + 6=521
521 * 6 + 5=3131
Here the sequence is first multiply by 2 and then add 1 AND
in the next number the reverse is following .
 
Shivansh Baranwal
2 years ago
1,3,7,25,103,......
The series can be represented as
1,1+2,3+4,12+13,52+51,260+261
We observe
Third term is second term higher value i.e.(2*1)*2+2*2-1
Fourth term is higher value 4*3+4*3+1
Fifth term is higher value 13*4+13*4-1
Sixth term will be higher value i.e. 52*5+52*5(+or -)1